OpenStudy (anonymous):

Please help! i have solution and answer but i still dont understand how they got to the answer.

OpenStudy (anonymous):

\[\int\limits_{}^{}\sqrt{2x+1}dx\]

OpenStudy (anonymous):

let u=2x+1 du=2dx dx=1/2 \[\int\limits_{}^{}\sqrt{u}*1/2du=1/2\int\limits_{}^{}u^{1/2}du \] \[=1/2*u^{3/2}/(3/2)+C\] =1/3(2x+1)6{3/2}+C

OpenStudy (anonymous):

why its sqrt u?? in the dirst place

OpenStudy (anonymous):

If u=2x +1, then the expression under the radical is u. So, sq rt of u is the same thing as u^(1/2).

OpenStudy (anonymous):

is it beacuse sqrt(2x+1)?

OpenStudy (anonymous):

kapeesh?

OpenStudy (anonymous):

whats that ?

OpenStudy (anonymous):

lol....too clear or not clear enough?

OpenStudy (anonymous):

@Hoa, no problem, any help from friends is always appreciated.

OpenStudy (anonymous):

teamwork play ! anyway, so why its sqrt (u)?

OpenStudy (anonymous):

then it says 2/2 du.

OpenStudy (anonymous):

1/2

OpenStudy (anonymous):

Sq rt is the same thing as the power 1/2

OpenStudy (anonymous):

I see. then = 1/2 in the front of sign of S u1/2 du. where did sqrt u go?

OpenStudy (anonymous):

Since sq rt is the same thing as 1/2, they change \[\int\limits_{}^{}\sqrt{u}du=\int\limits_{}^{}u ^{\frac{ 1 }{ 2 }}du\]

OpenStudy (anonymous):

The 1/2 in front of the integral is because u = 2x +1 and du=2dx, which means that xdx =(1/2)du. So the 1/2 just goes out in front of the integral.

OpenStudy (anonymous):

I see. these questions are really confusing. i will open open up the new question. would you help me with that?

OpenStudy (anonymous):

I have time for one more question. post it.

OpenStudy (anonymous):

thank you so much! k